# Meta-Computing Enhanced Federated Learning in IIoT: Satisfaction-Aware Incentive Scheme via DRL-Based Stackelberg Game

### v2

@@ Abstract
The Industrial Internet of Things (IIoT) leverages Federated Learning (FL) for distributed model training while preserving data privacy, and meta-computing enhances FL by optimizing and integrating distributed computing resources, improving efficiency and scalability. Efficient IIoT operations require a trade-off between model quality and training latency. Consequently, a primary challenge of FL in IIoT is to optimize overall system performance by balancing model quality and training latency. This paper designs a satisfaction function that accounts for data size, Age of Information (AoI), and training latency for meta-computing. Additionally, the satisfaction function is incorporated into the utility functions to incentivize nodes in IIoT participation in model training. We model the utility functions of servers and nodes as a two-stage Stackelberg game and employ a deep reinforcement learning approach to learn the Stackelberg equilibrium. This approach ensures balanced rewards and enhances the applicability of the incentive scheme for IIoT. Simulation results demonstrate that, under the same budget constraints, the proposed incentive scheme improves utility by at least 23.7% compared to existing FL schemes without compromising model accuracy.
@@ I Introduction
The Industrial Internet of Things (IIoT) serves as a core pillar of intelligent manufacturing and digital industrial systems, enabling real-time data collection, analysis, and intelligent decision making by connecting a large number of industrial devices, sensors, and intelligent systems [1, 2, 3]. Meta-computing [4] is an emerging computing paradigm that integrates and optimizes distributed computing resources in IIoT through key modules such as zero-trust computing management, task management, device management, identity and access management, and resource scheduling. These modules enable secure aggregation, dynamic task allocation, and efficient resource orchestration, providing efficient, flexible, and scalable computing services. Federated Learning (FL) [5, 6] is a collaborative AI approach that enables IIoT to optimize global machine learning models without migrating local data from IIoT nodes, thus enhancing the intelligent analytics capabilities of meta-computing while ensuring data privacy [7, 8].
@@
To ensure efficient operation of IIoT, FL in meta-computing must meet key requirements such as efficient resource allocation, high-quality models, and low latency [9]. However, due to the large scale of IIoT devices and the complexity of data structures, FL in meta-computing still faces several challenges.
@@
First, IIoT systems consist of numerous heterogeneous devices with varying computational capabilities and communication bandwidths, while task demands fluctuate dynamically with changes in the production environment. Some low-power devices struggle with high-load computational tasks, whereas high-performance computing nodes may experience bottlenecks due to unbalanced resource allocation, reducing overall system efficiency. Therefore, a key challenge is how to efficiently schedule tasks within limited resources to optimize resource allocation. Second, machine learning models are widely used in IIoT for intelligent prediction, anomaly detection, and automated control. However, high-quality models typically require significant computational resources and longer training times, while IIoT scenarios with large data volumes and complex structures often demand low latency and real-time responsiveness. Moreover, since IIoT nodes operate in distributed and autonomous environments, they are generally unwilling to share private information, further complicating the balance between model quality and latency [9]. Consequently, reducing computational and communication latency while maintaining model quality remains a critical research challenge.
@@
To address these challenges, we propose a novel evaluation metric, Satisfaction, to quantify the contribution of IIoT nodes in the FL process. This metric considers data volume, Age of Information (AoI) [10], and computation latency to assess node value. Specifically, AoI measures model quality, prioritizing newer models in updates, while computation latency evaluates node responsiveness, preventing high-latency devices from degrading overall training efficiency. Building on this metric, we develop a Satisfaction-Aware Incentive Mechanism that constructs utility functions for both nodes and the server, dynamically adjusts budget allocations, and incentivizes resource sharing. This mechanism encourages high-data-quality and high-efficiency nodes to actively participate in training. Furthermore, we model the incentive mechanism as a Stackelberg Game, where IIoT nodes act as followers responding to the server’s leader decisions. Unlike traditional game-theoretic approaches, we employ Deep Reinforcement Learning (DRL) to learn optimal strategies from historical experience, eliminating the need for private information from participating agents. Our main contributions are summarized as follows.
@@
  • We transform the optimization problem involving two utility functions into a Stackelberg Game Model and prove the uniqueness of the Stackelberg Equilibrium (SE). We then employ a DRL-based algorithm to learn the SE without requiring private information from other agents, enabling optimal strategy learning purely from historical experience.
@@ II-A FL in IIoT
FL has been widely applied in IIoT to enable decentralized intelligence while preserving data privacy. However, the heterogeneous nature of IIoT devices poses significant challenges for efficient resource scheduling. The authors in [11] proposed a semi-decentralized federated edge learning (SD-FEEL) framework that optimizes device association, resource allocation, and edge server placement to minimize training loss and improve accuracy within a limited cost budget. The authors in [12] proposed a client scheduling scheme for a semi-decentralized FL (SD-FL) framework in IIoT, optimizing client-server association to enhance communication reliability and training efficiency while minimizing global training loss. The authors in [13] proposed a FL-based Cyber Threat Intelligence framework (FL-CTIF) that enhances IIoT security by leveraging information fusion to improve cyberattack detection accuracy while reducing training rounds and CPU consumption. The authors in [14] proposed a decentralized federated learning framework for edge-enabled smart manufacturing, optimizing industrial data processing with inexact ADMM algorithms to enhance low-latency and secure decision-making while improving response time and maintaining accuracy. The authors in [15] proposed a DT-empowered IIoT architecture optimized with FL, integrating deep reinforcement learning for device selection and an asynchronous FL scheme to address device heterogeneity. The authors in [16] proposed a framework FDEI integrating federated learning with DT-enabled IIoT to enhance service quality and trustworthiness. The authors in [17] proposed an energy-efficient federated learning framework for digital twin-enabled IIoT, where IIoT devices dynamically select training methods based on environmental conditions. The authors in [18] introduced a meta-computing-driven vertical FL algorithm to address data incompleteness in heterogeneous IIoT devices, enabling collaborative intelligence without requiring shared feature spaces.
However, existing studies mainly focus on integrating DT with FL while lacking effective mechanisms for dynamic resource scheduling.
@@ II-B Incentive Schemes in IIoT
FL relies on distributed nodes to jointly participate in model training, and the performance of nodes directly affects the training latency and model quality [19]. Without an effective incentive scheme, selfish nodes might not provide sufficient resources or may be reluctant to join the federated learning task [20]. Therefore, designing an incentive scheme that fairly evaluates node contributions and encourages high-quality participation is crucial. The authors in [21] proposed the Model Value Transfer Incentive (MVTI) to enhance federated learning incentives for AIoT, ensuring secure execution and fair benefit redistribution with smart contracts and IPFS. The authors in [22] proposed BIMFL, a bilevel optimization-based incentive mechanism for federated learning, where the parameter server assigns rewards, and devices decide participation to maximize profit based on rewards and energy costs. The authors in [23] proposed a communication-efficient incentive mechanism for federated learning using Nash bargaining theory, formulating a concurrent bargaining game with a probabilistic greedy client selection algorithm. The authors in [24] proposed a multi-exit-based federated edge learning (ME-FEEL) framework for resource-constrained IIoT, enabling devices with limited computational power to exit early during training, reducing latency and maximizing participation in model aggregation. The authors in [25] proposed a social credit-based incentive mechanism to enhance cooperative D2D transmission across IIoT networks, improving throughput and reducing power consumption.
However, existing incentive mechanisms primarily focus on reward distribution and participation decisions but lack a comprehensive strategy to balance training latency and model quality. Most approaches fail to incorporate real-time factors such as data freshness and computational efficiency when evaluating node contributions, leading to suboptimal training performance in dynamic IIoT environments.
@@ II-C Game Theoretic and DRL Approaches in IIoT
Federated learning in IIoT involves complex resource allocation challenges. Traditional heuristic-based scheduling methods often fail to adapt to dynamic environments. Game-theoretic models have been explored to address resource competition in hierarchical decision-making scenarios, while DRL has emerged as a promising technique for learning optimal resource allocation policies in complex, dynamic IIoT networks. The authors in [26] proposed a blockchain-based resource trading framework for multi-UAV-assisted IIoT networks, leveraging Stackelberg game modeling and multi-agent DRL to optimize dynamic resource allocation while ensuring security and privacy. The authors in [27] proposed a DRL-based cloud-edge task scheduling model LsiA3CS, leveraging a Markov game and an asynchronous advantage actor-critic (A3C) algorithm to balance computing resources and reduce communication latency in large-scale IIoT networks. The authors in [28] focused on Stackelberg game-driven dynamic pricing and resource allocation, utilizing deep Q-networks (DQN) to approximate optimal strategies for resource pricing and task scheduling in IIoT.
@@ III System Model
In order to promote efficient task processing and decision making in IIoT, we design MEFL, a meta-computing enhanced FL framework, as shown in Fig. 1. The device management module contains multiple edge nodes, whose primary purpose is to collect data from production devices, integrate the computational, storage, and communication resources of the edge nodes, and map these resources to the server. These resources are then converted into objects that can be easily accessed by the resource scheduler, enabling local FL training based on the incentive scheme developed by the task manager. The resource scheduler module contains several digital twins of edge nodes, which continuously monitor changes in the configuration details of physical nodes, simulate their possible states, and dynamically optimize resource allocation. The task management module, located on the server, accepts user requests, decomposes tasks, and designs incentive schemes based on conditional constraints. The zero-trust computing management module performs global aggregation for FL via blockchain, while the identity and access management module ensures users have appropriate permissions to access data. Since the zero-trust computing management and identity and access management modules are not the focus of this paper, detailed discussions are omitted. Comprehensive analyses of these modules are available in [4]. The main notations used in this paper are shown in Table I.
We represent the set of IIoT nodes involved as ${I_{all}}=\{1,...,i,...,I\}$ and the task duration as $T$. Upon task arrival, each model training is iterated $K$ times to minimize global losses, where $K$ is specified by the server. Assume that there are $I$ nodes with local data sets $\{D^{\prime}_{1},D^{\prime}_{2},...,D^{\prime}_{I}\}$. We define $D_{i}\triangleq|D^{\prime}_{i}|$, where ||$\cdot$|| ⋅ | denotes the size of the data set with meta-computing, each node downloads a shared global model $\omega$ from the server and trains the model using its local data. The node then uploads the new local model update to the server. Therefore, the total size of data samples from $I$ nodes is $\sum\nolimits_{i=1}^{I}{{D}_{i}={D}_{all}}$. The loss function of the node $i$ with the data set $D^{\prime}_{i}$ is
@@
Due to the inherent complexity of many machine learning models, it is hard to find a closed-form solution. Therefore, it is often solved using gradient descent techniques [29]. Given that IIoT nodes may be reluctant to provide new sensing data for time-sensitive FL tasks, reliable incentives are greatly needed to encourage users to share new sensing data, which is discussed in the next section.
@@
$Q_{i}$ The quality of the model provided by node $i$
@@ IV Satisfaction: Quality Control for IIoT
The structure of IIoT is complex, containing multiple types of industrial devices and involving a large number of industrial nodes, where the trade-off between low latency and high model quality must be carefully considered in meta-computing enhanced federated learning. Latency plays a critical role in industrial control and decision-making processes. High latency can lead to delayed data synchronization, causing inefficiencies in predictive maintenance, anomaly detection, and real-time monitoring [30]. Meanwhile, model quality determines the accuracy and reliability of decision-making. Degraded model quality means that the data used for decision-making is less accurate or outdated, which may lead to incorrect control actions, reduced production efficiency, or potential safety risks. Furthermore, model quality directly influences the effectiveness of IIoT decision-making processes—low-quality models may lead to suboptimal system adjustments or improper responses to changing industrial conditions. Traditional quality metrics in FL do not adequately capture the requirements of IIoT environments. Therefore, we propose a satisfaction metric $G_{i}$, to balance quality and latency. The satisfaction metric $G_{i}$ of the task $i$ is denoted as
$\begin{split}{G_{i}}=\tau{Q_{i}}-\lambda{E_{i}},\end{split}$ (4)
where $\tau$ and $\lambda$ are the conversion parameters for quality and latency, respectively.
Since the timeout of gradient updates can negatively impact the learning results, we use AoI to measure the freshness of the model in order to ensure its quality. AoI, as a valid measure of information freshness, denotes the latency of the information from its generation to the completion of the model training after it has been uploaded to the server [31], and it can enhance the performance of time-critical applications and services. In FL, we assume that requests arrive at the beginning of each cycle. We focus on FL with a data cache buffer on the node and AoI [32]. The node $i$ periodically updates its cached data with an update period $\theta_{i}$ independently. It is denoted as
@@
where ${c_{i}}t({c_{i}}\in N)$ is the time spent by the node $i$ to collect and process the model training data; ${a_{i}}t({a_{i}}\in N)$ is the duration from the end of data collection to the beginning of the next phase of data collection. It may contains service time period and idle time period.
When the request arrives during the data collection phase or at the beginning of phase $(c_{i}+1)t$, the AoI is $t$. This is the minimum AoI value. For requests arriving in phase $l\times t$, where $l\geq(c_{i}+2)$, the AoI will be $[l-(c_{i}+1)+1]t$. We suppose that $t$ is fixed and that the update cycle ${\theta_{i}}$ is affected by ${c_{i}}$ and ${a_{i}}$. Let us consider a case with an adjustable update phase, i.e., when ${a_{i}}=a$ is fixed, ${c_{i}}=\frac{{{\theta_{i}}}}{t}-a$, we use ${\theta_{i}}$ instead of ${c_{i}}$. Therefore, the average AoI of the node $i$ is
$\begin{split}{{\bar{A}}_{i}}&=\frac{t}{{{c_{i}}+{a_{i}}}}[{c_{i}}+1+\frac{{({a_{i}}-1)({a_{i}}+2)}}{2}]\\ &=\frac{{t{\theta_{i}}}}{{{\theta_{i}}-at}}+\frac{{{t^{2}}}}{{{\theta_{i}}-at}}(\frac{{{a^{2}}-a}}{2}),\end{split}$ (6)
@@
In addition to AoI, traditional FL experiences service latency [33], defined in this paper as ${E_{i}}$. Unlike AoI, service latency refers to the duration from when the node receives a request until it uploads a local model. It includes both the data collection period and the model training period. The probability of a request arriving is uniformly distributed across periods, i.e., $\frac{1}{T}$. If the request arrives in the $n$th period of the data collection cycle, the service delay is ${c_{i}}t+t-(n-1)t$. If the request arrives in any of the remaining periods, the service latency is $t$. Therefore, the average service latency is
$\begin{split}{{\bar{E}}_{i}}&=\frac{{{c_{i}}}}{{{c_{i}}+{a_{i}}}}[({c_{i}}t+t)+\cdots+({c_{i}}t+t-({c_{i}}-1)t)]+\frac{{{a_{i}}}}{{{c_{i}}+{a_{i}}}}t\\ &=\frac{{{c_{i}}}}{{{c_{i}}+{a_{i}}}}[\frac{{{c_{i}}}}{2}({c_{i}}+3)]+\frac{{{a_{i}}}}{{{c_{i}}+{a_{i}}}}t\\ &=\frac{{{{({\theta_{i}}-at)}^{3}}}}{{2t{\theta_{i}}}}+\frac{{3{{({\theta_{i}}-at)}^{2}}}}{{2{\theta_{i}}}}+\frac{{a{t^{2}}}}{\theta}.\end{split}$ (7)
@@
$\begin{split}{Q_{i}}=\rho\frac{{{D_{i}}}}{{{A_{i}}}}=\rho\frac{{Td({\theta_{i}}-at)}}{{{\theta_{i}}(t{\theta_{i}}+{t^{2}}(\frac{{{a^{2}}-a}}{2}))}},\end{split}$ (8)
where $\rho$ is the parameter used to adjust the quality.

### v3

@@ Abstract
The Industrial Internet of Things (IIoT) leverages Federated Learning (FL) for distributed model training while preserving data privacy, and meta-computing enhances FL by optimizing and integrating distributed computing resources, improving efficiency and scalability. Efficient IIoT operations require a trade-off between model quality and training latency. Consequently, a primary challenge of FL in IIoT is to optimize overall system performance by balancing model quality and training latency. This paper designs a satisfaction function that accounts for data size, Age of Information (AoI), and training latency for meta-computing. Additionally, the satisfaction function is incorporated into the utility function to incentivize IIoT nodes to participate in model training. We model the utility functions of servers and nodes as a two-stage Stackelberg game and employ a deep reinforcement learning approach to learn the Stackelberg equilibrium. This approach ensures balanced rewards and enhances the applicability of the incentive scheme for IIoT. Simulation results demonstrate that, under the same budget constraints, the proposed incentive scheme improves utility by at least 23.7% compared to existing FL schemes without compromising model accuracy.
@@ I Introduction
The Industrial Internet of Things (IIoT) serves as a core pillar of intelligent manufacturing and digital industrial systems, enabling real-time data collection, analysis, and intelligent decision making by connecting a large number of industrial devices, sensors, and intelligent systems [c1, c2, c3]. Meta-computing [b47] is an emerging computing paradigm that integrates and optimizes distributed computing resources in IIoT through key modules such as zero-trust computing management, task management, device management, identity and access management, and resource scheduling. These modules enable secure aggregation, dynamic task allocation, and efficient resource orchestration, thereby providing flexible and scalable computing services. Federated Learning (FL) [b8, b9] is a collaborative AI approach that enables IIoT to optimize global machine learning models without migrating local data from IIoT nodes, thus enhancing the intelligent analytics capabilities of meta-computing while ensuring data privacy [b10, b50].
@@
To ensure efficient operation of IIoT, FL in meta-computing must meet key requirements such as efficient resource allocation, high-quality models, and low latency [c6]. However, due to the large scale of IIoT devices and the complexity of data structures, FL in meta-computing still faces several challenges.
@@
First, IIoT systems consist of numerous heterogeneous devices with varying computational capabilities and communication bandwidths, while task demands fluctuate dynamically with changes in the production environment. Some lowpower devices struggle with high-load computational tasks, whereas high-performance computing nodes may experience bottlenecks due to unbalanced resource allocation, reducing overall system efficiency [c25]. Therefore, a key challenge is how to efficiently schedule tasks within limited resources to optimize resource allocation. Second, machine learning models are widely used in IIoT for intelligent prediction, anomaly detection, and automated control. However, high-quality models typically require significant computational resources and longer training times, while IIoT scenarios with large data volumes and complex structures often demand low latency and real-time responsiveness. Moreover, since IIoT nodes operate in distributed and autonomous environments, they are generally unwilling to share private information, further complicating the balance between model quality and latency [c6]. Consequently, reducing computational and communication latency while maintaining model quality remains a critical research challenge.
@@
The problem addressed in this paper is how to design an efficient incentive scheme that enables distributed IIoT devices with heterogeneous resources to collaboratively train high-quality models while minimizing communication and computation latency under limited resource and privacy constraints. We propose a novel evaluation metric, Satisfaction, to quantify the contribution of IIoT nodes in the FL process. This metric considers data volume, Age of Information (AoI) [b40], and computation latency to assess node value. Specifically, AoI reflects the freshness of model updates, prioritizing newer information, while computation latency evaluates node responsiveness, preventing high-latency devices from degrading overall training efficiency. Building on this metric, we develop a satisfaction-aware incentive mechanism that jointly formulates the utility functions of the server and participating nodes, dynamically allocates budgets, and incentivizes efficient resource sharing. This mechanism encourages nodes with high data quality and computational efficiency to actively participate in model training. Furthermore, the incentive mechanism is modeled as a Stackelberg Game, where IIoT nodes act as followers responding to the server’s leader decisions. Unlike traditional game-theoretic approaches, we employ Deep Reinforcement Learning (DRL) to learn optimal strategies through interaction experience, thereby eliminating the need for participating agents’ private information. The main contributions of this paper are summarized as follows:
@@
  • We transform the optimization problem involving two utility functions into a Stackelberg game model and prove the uniqueness of the Stackelberg Equilibrium (SE). A DRL-based algorithm is employed to learn the SE without requiring private information from other agents, enabling optimal strategy learning purely from historical experience.
@@ II-A FL in IIoT
FL has been widely applied in IIoT to enable decentralized intelligence while preserving data privacy. However, the heterogeneous nature of IIoT devices poses significant challenges for efficient resource scheduling. The authors in [c19] proposed a semi-decentralized federated edge learning (SD-FEEL) framework that optimizes device association, resource allocation, and edge server placement to minimize training loss and improve accuracy within a limited cost budget. The authors in [c20] proposed a client scheduling scheme for a semi-decentralized FL (SD-FL) framework in IIoT, optimizing client-server association to improve communication reliability and training efficiency while minimizing global training loss. The authors in [c21] proposed a FL-based Cyber Threat Intelligence framework (FL-CTIF) that improves IIoT security by using information fusion to improve cyberattack detection accuracy while reducing training rounds and CPU consumption. The authors in [c22] proposed a decentralized federated learning framework for edge-enabled smart manufacturing, optimizing industrial data processing with inexact ADMM algorithms to improve low-latency and secure decision-making while improving response time and maintaining accuracy. The authors in [c7] proposed a DT-empowered IIoT architecture optimized with FL, integrating deep reinforcement learning for device selection and an asynchronous FL scheme to address device heterogeneity. The authors in [c8] proposed a framework FDEI integrating federated learning with DT-enabled IIoT to enhance service quality and trustworthiness. The authors in [c9] proposed an energy-efficient federated learning framework for digital twin-enabled IIoT, where IIoT devices dynamically select training methods based on environmental conditions. The authors in [c24] introduced a meta-computing driven vertical FL algorithm to address data incompleteness in heterogeneous IIoT devices, enabling collaborative intelligence without requiring shared feature spaces.
However, existing studies still struggle to efficiently schedule tasks in IIoT scenarios where heterogeneous device resources are limited and dynamically varying. The proposed MEFL framework leverages meta-computing for resource scheduling to optimize resource allocation and incorporates a satisfaction-aware incentive scheme, thereby enabling intelligent decision-making and improved resource utilization in IIoT environments.
@@ II-B Incentive Schemes in IIoT
FL relies on distributed nodes to jointly participate in model training, and the performance of nodes directly affects the training latency and model quality [c10]. Without an effective incentive scheme, selfish nodes might not provide sufficient resources or may be reluctant to join the federated learning task [b20]. Therefore, designing an incentive scheme that fairly evaluates node contributions and encourages high-quality participation is crucial. The authors in [c11] proposed the Model Value Transfer Incentive (MVTI) to enhance federated learning incentives for AIoT, ensuring secure execution and fair benefit redistribution with smart contracts and IPFS. The authors in [c12] proposed BIMFL, a bilevel optimization-based incentive mechanism for federated learning, where the parameter server assigns rewards, and devices decide participation to maximize profit based on rewards and energy costs. The authors in [c13] proposed a communication-efficient incentive mechanism for federated learning using Nash bargaining theory, formulating a concurrent bargaining game with a probabilistic greedy client selection algorithm. The authors in [c18] proposed a multi-exit-based federated edge learning (ME-FEEL) framework for resource-constrained IIoT, enabling devices with limited computational power to exit early during training, reducing latency and maximizing participation in model aggregation. The authors in [c23] proposed a social credit-based incentive mechanism to enhance cooperative D2D transmission across IIoT networks, improving throughput and reducing power consumption.
Although the above studies improve the fairness, participation, and communication efficiency of FL, they still struggle to simultaneously meet latency constraints and ensure model quality in dynamic IIoT scenarios. The proposed satisfaction-aware incentive scheme jointly considers data size, AoI, and service latency to quantify node contributions and balance model quality against service latency, thereby achieving an efficiency–accuracy trade-off in IIoT.
@@ II-C Game Theoretic and DRL Approaches in IIoT
Federated learning in IIoT involves complex resource allocation challenges. Traditional heuristic-based scheduling methods often fail to adapt to dynamic environments. Game-theoretic models have been explored to address resource competition in hierarchical decision-making scenarios, while DRL has emerged as a promising technique for learning optimal resource allocation policies in complex, dynamic IIoT networks. The authors in [c14] proposed a blockchain-based resource trading framework for multi-UAV-assisted IIoT networks, leveraging Stackelberg game modeling and multi-agent DRL to optimize dynamic resource allocation while ensuring security and privacy. The authors in [c15] proposed a DRL-based cloud-edge task scheduling model LsiA3CS, leveraging a Markov game and an asynchronous advantage actor-critic (A3C) algorithm to balance computing resources and reduce communication latency in large-scale IIoT networks. The authors in [c16] focused on Stackelberg game-driven dynamic pricing and resource allocation, utilizing deep Q-networks (DQN) to approximate optimal strategies for resource pricing and task scheduling in IIoT.
@@ III System Model
In order to promote efficient task processing and decision making in IIoT, we design MEFL, a meta-computing enhanced FL framework, as shown in Fig. 1. The device management module contains multiple edge nodes, whose primary purpose is to collect data from production devices, integrate the computational, storage, and communication resources of the edge nodes, and map these resources to the server. These resources are then converted into objects that can be easily accessed by the resource scheduler, enabling local FL training based on the incentive scheme developed by the task manager. The resource scheduler module contains several digital twins of edge nodes, which continuously monitor changes in the configuration details of physical nodes, simulate their possible states, and dynamically optimize resource allocation. The task management module, located on the server, accepts user requests, decomposes tasks, and designs incentive schemes based on conditional constraints. The zero-trust computing management module performs global aggregation for FL via blockchain, while the identity and access management module ensures users have appropriate permissions to access data. Since the zero-trust computing management and identity and access management modules are not the focus of this paper, detailed discussions are omitted. Comprehensive analyses of these modules are available in [b47].
In MEFL, users view devices on a virtual platform and publish tasks, which are transmitted to the task management module on server. The server interacts with virtual nodes in the resource scheduling module via a Stackelberg game and generates a satisfaction-aware incentive scheme based on the game results. The incentives are distributed to physical nodes in the device management module to drive local FL tasks. After completing model training, the physical nodes upload their updates to the zero-trust computing management module for aggregation. After multiple iterations, the final global model is produced and returned to the user, while the user completes the corresponding incentive payment. The main notations used in this paper are shown in Table I.
We represent the set of IIoT nodes involved as ${I_{all}}=\{1,...,i,...,I\}$ and the task duration as $T$. Upon task arrival, each model training is iterated $K$ times to minimize global losses, where $K$ is specified by the server. Assume that there are $I$ nodes with local data sets $\{D^{\prime}_{1},D^{\prime}_{2},...,D^{\prime}_{I}\}$. We define $D_{i}\triangleq|D^{\prime}_{i}|$, where $|\cdot|$ denotes the size of the data set with meta-computing, each node downloads a shared global model $\omega$ from the server and trains the model using its local data. The node then uploads the new local model update to the server. Therefore, the total size of data samples from $I$ nodes is $\sum\nolimits_{i=1}^{I}{{D}_{i}={D}_{all}}$. The loss function of the node $i$ with the data set $D^{\prime}_{i}$ is
@@
Meta-computing assists the training process and the exchange of models between nodes and the server. Specifically, it accelerates local update procedures and improves the efficiency of parameter transmission. As a result, each training round can complete sooner, and nodes can submit updates more quickly. This architectural support enhances the timeliness of federated learning and reduces resource overhead.
Due to the inherent complexity of many machine learning models, it is hard to find a closed-form solution. Therefore, it is often solved using gradient descent techniques [b37]. Given that IIoT nodes may be reluctant to provide fresh sensing data for time-sensitive FL tasks, reliable incentives are greatly needed to encourage users to share fresh sensing data, which is discussed in the next section.
@@
$M_{i}$ The quality of the model provided by node $i$
@@ IV Satisfaction: Quality Control for IIoT
The structure of IIoT is complex, containing multiple types of industrial devices and involving a large number of industrial nodes, where the trade-off between low latency and high model quality must be carefully considered in meta-computing enhanced federated learning. Latency plays a critical role in industrial control and decision-making processes. High latency can lead to delayed data synchronization, causing inefficiencies in predictive maintenance, anomaly detection, and real-time monitoring [c17]. Meanwhile, model quality determines the accuracy and reliability of decision-making. Degraded model quality means that the data used for decision-making is less accurate or outdated, which may lead to incorrect control actions, reduced production efficiency, or potential safety risks. Furthermore, model quality directly influences the effectiveness of IIoT decision-making processes—low-quality models may lead to suboptimal system adjustments or improper responses to changing industrial conditions. Traditional quality metrics in FL do not adequately capture the requirements of IIoT environments. Therefore, we propose a satisfaction metric $G_{i}$, to balance quality and latency. The satisfaction metric $G_{i}$ of the task $i$ is denoted as
$\begin{split}{G_{i}}=\tau{M_{i}}-\lambda{E_{i}},\end{split}$ (4)
where $\tau>0$ and $\lambda>0$ are weighting parameters. $\tau$ weights the contribution of model quality to satisfaction, while $\lambda$ weights the penalty imposed by service latency.
Since the timeout of gradient updates can negatively impact the learning results, we use AoI to measure the freshness of the model in order to ensure its quality. AoI, as a valid measure of information freshness, denotes the latency of the information from its generation to the completion of the model training after it has been uploaded to the server [b35], and it can enhance the performance of time-critical applications and services. In FL, we assume that requests arrive at the beginning of each cycle. We focus on FL with a data cache buffer on the node and AoI [b38]. The node $i$ periodically updates its cached data with an update period $\theta_{i}$ independently. It is denoted as
@@
where ${c_{i}}t\ ({c_{i}}\in N)$ is the data collection and preprocessing time at node $i$ with meta-computing assistance, and ${a_{i}}t\ ({a_{i}}\in N)$ is the duration from the end of data collection to the beginning of the next phase of data collection, which includes the meta-computing assisted service period and any idle time.
When the request arrives during the data collection phase or at the beginning of phase $(c_{i}+1)t$, the AoI is $t$. This is the minimum AoI value. For requests arriving in phase $l\times t$, where $l\geq(c_{i}+2)$, the AoI will be $l-(c_{i}+1)+1]t$. We suppose that $t$ is fixed and that the update cycle $\theta_{i}$} is affected by $c_{i}$} and $a_{i}$}. Let us consider a case with an adjustable update phase, i.e., when ${a_{i}}=a$ is fixed, ${c_{i}}=\frac{{{\theta_{i}}}}{t}-a$, we use $\theta_{i}$} instead of $c_{i}$}. Therefore, the average AoI of the node $i$ is
$\begin{split}{A_{i}}&=\frac{t}{{{c_{i}}+{a_{i}}}}[{c_{i}}+1+\frac{{({a_{i}}-1)({a_{i}}+2)}}{2}]\\ &=\frac{{t{\theta_{i}}}}{{{\theta_{i}}-at}}+\frac{{{t^{2}}}}{{{\theta_{i}}-at}}(\frac{{{a^{2}}-a}}{2}),\end{split}$ (6)
@@
In addition to AoI, traditional FL experiences service latency [b39], defined in this paper as $E_{i}$}. Unlike AoI, service latency refers to the duration from when the node receives a request until it uploads a local model. It includes both the data collection period and the model training period. The probability of a request arriving is uniformly distributed across periods, i.e., $\frac{1}{T}$. If the request arrives in the $n$th period of the data collection cycle, the service delay is ${c_{i}}t+t-(n-1)t$. If the request arrives in any of the remaining periods, the service latency is $t$. Therefore, the average service latency is
$\begin{split}{E_{i}}&=\frac{{{c_{i}}}}{{{c_{i}}+{a_{i}}}}[({c_{i}}t+t)+\cdots+({c_{i}}t+t-({c_{i}}-1)t)]+\frac{{{a_{i}}}}{{{c_{i}}+{a_{i}}}}t\\ &=\frac{{{c_{i}}}}{{{c_{i}}+{a_{i}}}}[\frac{{{c_{i}}}}{2}({c_{i}}+3)]+\frac{{{a_{i}}}}{{{c_{i}}+{a_{i}}}}t\\ &=\frac{{{{({\theta_{i}}-at)}^{3}}}}{{2t{\theta_{i}}}}+\frac{{3{{({\theta_{i}}-at)}^{2}}}}{{2{\theta_{i}}}}+\frac{{a{t^{2}}}}{\theta}.\end{split}$ (7)
@@
$\begin{split}{M_{i}}=\rho\frac{{{D_{i}}}}{{{A_{i}}}}=\rho\frac{{Td({\theta_{i}}-at)}}{{{\theta_{i}}(t{\theta_{i}}+{t^{2}}(\frac{{{a^{2}}-a}}{2}))}},\end{split}$ (8)
where $\rho>0$ is a quality scaling factor that transforms the AoI based ratio into a normalized quality measure, ensuring comparability with the latency penalty term.